# How Transformers Solve Propositional Logic Problems: A Mechanistic Analysis

### v2

@@ B.2.2 Training details
B.3 Answer for the LogOp Chain
@@
Evidence 3c: identification of LogOp-chain starting nodes at layer 3. Attention heads (3,1) and (3,3), when concatenated, produce embeddings which we can linearly decode the two starting nodes of the LogOp chain with test accuracy greater than 98%. We also find that they focus their attention in the rule section of the context (as shown in Figure 8). Due to causal attention, this means that they determine the two starting nodes from the LogOp-relevant rules.
@@
Evidence 3d: layer-2 residual stream at ANSWER is important to correct prediction. We verify that layer-3 attention does rely on information in the layer-2 residual stream (at the ANSWER position):
@@
Figure 8: Attention statistics, averaged over 500 samples, all of which query for the LogOp chain. The x-axis is simply an example prompt that helps illustrate where the attention is really placed at. Observe that only attention head (3,2) pays significant attention to the fact section. The other two heads focus on the rule section.
@@ B.4 Extra remarks
Appendix C Mistral-7B experiment details
@@ C.1 Problem format
C.2 Finer details of query-based activation patching
@@
C.2.1 QUERY-based activation patching experiments: metrics
@@
$\frac{\frac{1}{N}\sum_{n\in[N]}\Delta_{orig\to alt,n;(\ell,h,t)}-\Delta_{orig}^{\dagger}}{\Delta_{alt}^{\dagger}-\Delta_{orig}^{\dagger}}.$ (1)
where $\Delta_{orig}^{\dagger}=\frac{1}{N}\sum_{n\in[N]}\text{logit}({\bm{X}}_{n})[y_{alt,n}]-\text{logit}({\bm{X}}_{n})[y_{orig,n}]$, and $\Delta_{alt}^{\dagger}=\frac{1}{N}\sum_{n\in[N]}\text{logit}({\bm{X}}^{\prime}_{n})[y_{alt,n}]-\text{logit}({\bm{X}}^{\prime}_{n})[y_{orig,n}]$. The closer to 1 this score is, the stronger the model’s “belief” is altered; the closer to 0 it is, the closer the model’s “belief” is to the original unaltered one.
@@
C.2.2 Attention patterns of QUERY-sensitive attention heads
@@
We can observe that head (12,9) has the “cleanest” attention pattern out of the ones identified, placing on average $93.0\pm 9.4\%$ of it attention on the correct conclusion variable alone. The more diluted attention patterns of the other heads likely contribute to their weaker intervened logit difference score shown in Section 4.3.1 in the main text.
@@
An extra note about head (16,0): it does not primarily act like a “mover” head, as its attention statistics suggest that it processes the mixture of information from the QUERY position and the “:” position. Therefore, while we present its statistics along with the other queried-rule mover heads here since it does allocate significant attention weight on the QUERY position on average, we do not list it as such in the circuit diagram of Figure 3.
@@
Figure 11: Average attention weights of the fact processing heads, along with the standard deviations. The weights are calculated by dividing the actual attention weight placed in the Fact section by the total amount of attention placed in the relevant context (i.e. the part of the prompt excluding the 6 in-context examples).
@@
An interesting side note worth pointing out is that, (17,25) tends to only concentrate its attention in the facts section, similar to the fact-processing heads. The reason which we do not classify it as a fact-processing head and instead as a decision head is that, in addition to finding that their attention patterns tend to concentrate on the correct fact, evidence presented in subsection C.4 below suggest that they are not directly responsible for locating and moving the facts information to the “:” position, while the heads (16,12;14) exhibit such tendency strongly.
C.3 Queried-rule location interventions: analyzing the queried-rule locating heads
In this experiment, we only swap the location of the linear rule with the LogOp rule in the Rule section of the question, while keeping everything else the same (including all the in-context examples); we choose to query for the linear chain in every sample. As an example, we alter “Rules: A or B implies C. D implies E.” to “Rules: D implies E. A or B implies C.” while keeping everything else the same. The two prompts have the same answer.
Figure 13: Key activations patching results. In this experiment, we swap the location of the linear rule and the LogOp rule in the Rule section and keep everything else in the prompt the same; we patch the key activations of the attention heads in the Rule section only. (a) visualizes the average increase in the cross-entropy loss with respect to the true target (the true first token of the answer) for all key indices, and (b) shows the average and standard deviation of the top three key indices with the highest loss increase. Observe that these are the keys for the queried-rule locating heads (12,9), (14,24;26) and (9,25;26) identified in Section 4.3.1.
@@ C.3 Queried-rule location interventions: analyzing the queried-rule locating heads
The average increase in cross-entropy loss exhibit a trend which corroborate the hypothesis above, shown in Figure 13. While the average cross-entropy loss on the original samples is 0.463, patching (12,9), (14,24;26) and (9,25;26)’s keys (with corresponding key indices (12,2), (14,6) and (9,6)) in the Rule section result in loss increase of $0.375\pm 0.246$, $0.148\pm 0.091$ and $0.260\pm 0.293$ respectively on average (81%, 32% and 56% increase). Patching the other QUERY-sensitive attention heads’ keys in the Rule section, in contrast, show significantly smaller influence on the loss on average, telling us that their responsibilities are much less involved with directly finding or locating the queried rule via attention.
@@
While attention heads with key index (15,5) (i.e. heads (15, 20-23)) did not exhibit nontrivial sensitivity to QUERY-based patching (discussed in Section 4.3.1 in the main text), patching this key activation does result in a nontrivial $0.101\pm 0.113$ increase in loss. Examining the attention heads belonging to this group, we find that they indeed also perform the function of locating the queried rule similar to head (12,9). We find them to be less accurate and place less attention on the exact rule being queried on average, however: this weaker “queried-rule locating ability” likely contributed to their low scores in the QUERY-based patching experiments presented in the main text.
C.4 Facts interventions: analyzing the fact-processing and decision heads

### v3

@@ B.2.2 Training details
B.3 High-level reasoning strategy of the 3-layer transformer
We complement the text description of the reasoning strategy of the 3-layer transformer in the main text with Figure 8 below. It not only presents the main strategy of the model, but also summarizes the core evidence for specific parts of the strategy.
Figure 8: High-level overview of how the 3-layer transformer solves the logic problem. As shown in the grey blocks on the left, the model performs “routing” in layer 2 by sending a routing signal ${\bm{h}}_{route}$ to layer 3 (with its “sign” dependent on the query type), then the layer-3 attention block acts according to the “sign” of the routing signal sent to it. The middle (right) chain shows the strategy when the problem queries for linear (LogOp) chain.
B.4 Answer for the LogOp Chain
@@
Evidence 3d: linear non-decodability of linear chain’s answer. To provide further contrasting evidence for the linear decodability of the LopOp chain’s answer, we experimentally show that it is not possible to linearly decode the answer of the linear chain in the model. Due to the causal nature of the reasoning problem (it is only possible to know the answer at or after the QUERY token position), and the causal nature of the decoder-only transformer, we train a set of linear classifiers on all token positions at or after the QUERY token and up to the ANSWER token, and on all layers of the residual stream of the transformer. We follow the same procedure as in Evidence 3c, except in this set of experiments, for contrasting evidence, QUERY is for the LopOp chain, while the classifier is trained to predict the answer of the Linear chain. The maximum test accuracy of the linear classifiers across all aforementioned token positions and layer indices is only 32.7%. Therefore, the answer of the Linear chain is not linearly encoded in the model when QUERY is for the LopOp chain.
Evidence 3e: layer-2 residual stream at ANSWER is important to correct prediction. We verify that layer-3 attention does rely on information in the layer-2 residual stream (at the ANSWER position):
@@
Figure 9: Attention statistics, averaged over 500 samples, all of which query for the LogOp chain. The x-axis is simply an example prompt that helps illustrate where the attention is really placed at. Observe that only attention head (3,2) pays significant attention to the fact section. The other two heads focus on the rule section. Note that none of them concentrate attention on the QUERY token.
@@ B.5 Extra remarks
Appendix C The reasoning circuit in Mistral-7B: experimental details
@@ C.1 Problem format
C.2 Causal mediation analysis: further explanations
This subsection complements the causal mediation analysis methodologies we presented in §4.2 in the main text. In particular, we aim to visualize how the interventions are done in the circuit discovery and verification processes, by using a 2-layer 2-head transformer as an example for simplicity.
Figure 10 illustrates the activation patching procedure of circuit discovery. Recall that we are studying how a component inside the transformer causally influences the output of the model. In the specific example, we show how we would examine the causal influence of attention head (0,2)’s activations on the correct inference of the model.
Figure 10: Illustration of how activation patching is performed in the circuit-discovery process (necessity-based patching). We use a 2-layer 2-head transformer as a simplified example here, and use $(\ell,h)$ to denote an attention head.
In this specific illustrated example, we are studying the causal influence of attention head (0,2)’s activation on the correct inference of the model. After caching the altered activations of (0,2) (shown on the left), we run the model on the original prompt and cache the activations (shown in the middle), then replace the original activation of head (0,2) by its altered activations, and let the rest of the layers be computed normally (shown on the right) — they now operate out of distribution, and are colored in green.
In this specific example, the intervened run outputs logits which reflect “belief altering”: that is, the probability for the answer token of the original prompt now is lower than the answer token for the altered prompt. This indicates that head (0,2) has causal influence on the corrent inference of the model.
Circuit verification, on the other hand, goes through a somewhat more complex process of interventions, as illustrated in Figure 11. Recall our main procedure (discussed in the main text).
  1. 1.
    We run the LLM on the original prompts, and cache the activations of the attention heads.
  2. 2.
    Now, we run the LLM on the corresponding altered prompts, however, we freeze all the attention heads’ activations inside the model to their activations on the original prompts, except for those in the circuit $\mathcal{C}$ which we wish to verify (i.e. only the attention heads in $\mathcal{C}$ are allowed to run normally). We record the (circuit-intervened) altered logit differences on the altered prompts.
  3. 3.
    We average the circuit-intervened altered logit differences across the samples, namely $\frac{1}{N}\sum_{n=1}^{N}\Delta_{orig\to alt}^{\mathcal{C}}$ [^9] This specific term reflects, on average, how much the model favors outputting the answer token for the altered prompts over the original prompts after the circuit interventions., and check whether they approach the “maximal” altered logit difference, namely $\Delta_{alt}$ [^10] Recall that this term is obtained by running the LLM on the altered prompts without any modification to its internal activations at all. This specific term reflects, on average, how much the (un-intervened) model favors outputting the answer tokens for the altered prompts over those of the original prompts, when it is run on the altered prompts..
Remark.
As the reader can observe, we do not freeze the MLPs in our intervention experiments. We note that the MLPs do not move information between the residual streams at different token positions, as they only perform processing of whatever information present at the residual stream. Therefore, similar to Wang et al. (2023), we consider the MLPs as part of the “direct” path between two attention heads, and allow information to flow freely through them, instead of freezing them and disrupting the information flow between attention heads.
Figure 11: Illustration of how activation patching is performed in the circuit-verification process (sufficiency-based patching). We use a 2-layer 2-head transformer as a simplified example here. We use $(\ell,h)$ to denote an attention head.
In this specific illustrated example, we are verifying whether the circuit $\mathcal{C}=\{(0,2),(1,1)\}$ consisting of the two attention heads is sufficient for altering the “belief” of the model.
We first obtain $\Delta_{alt}$, the average altered logit difference, by running the model on the altered prompts without interventions. We also run the model on the original prompts and cache the attention heads’ activations (these “original” activations are colored in blue in the figure).
The naive “baseline” for sufficiency verification is the null circuit $\mathcal{C}_{null}=\emptyset$ (shown in the middle): we freeze all the attention heads to their original activations when running the model on the altered prompts. This null circuit, as shown in this example, barely alters the model’s “belief” from the original, as $\Delta^{{\color[rgb]{0,0,1}\mathcal{C}_{null}}}_{orig\to alt}\approx-\Delta_{alt}$, i.e. the model still strongly favors outputting the answer tokens for the original prompts over those of the altered prompts on average.
In contrast, if we unfreeze the attention heads in the circuit $\mathcal{C}$ when running the model (shown on the right), we observe that the model’s circuit-intervened logit difference approaches the “maximal” altered logit difference $\Delta_{alt}$. This indicates that the attention heads in $\mathcal{C}$ are sufficient for correctly manipulating the information flow (and processing the information) for reaching the right answer.
C.3 Finer details of query-based activation patching
@@
C.3.1 QUERY-based activation patching experiments: metrics
@@
$\frac{\frac{1}{N}\sum_{n\in[N]}\Delta_{orig\to alt,n;(\ell,h,t)}-\Delta_{orig}^{\dagger}}{\Delta_{alt}-\Delta_{orig}^{\dagger}}.$ (9)
where $\Delta_{orig}^{\dagger}=\frac{1}{N}\sum_{n\in[N]}\text{logit}({\bm{X}}_{orig,n})[y_{alt,n}]-\text{logit}({\bm{X}}_{orig,n})[y_{orig,n}]$, and $\Delta_{alt}=\frac{1}{N}\sum_{n\in[N]}\text{logit}({\bm{X}}_{alt,n})[y_{alt,n}]-\text{logit}({\bm{X}}_{alt,n})[y_{orig,n}]$. The closer to 1 this score is, the stronger the model’s “belief” is altered; the closer to 0 it is, the closer the model’s “belief” is to the original unaltered one.
@@
C.3.2 Attention head group patching
Figure 12: Attention head patching, highlighting the ones with the highest intervened logit difference; $x$-axis is the head index. (a) shows single-head patching results (same as the one shown in the main text, repeated here for the reader’s convenience), and (b) shows a coarser-grained head patching in groups. In (b), we only highlight the head groups that are not captured well by (a).
We note that Grouped-Query Attention used by Mistral-7B adds subtlety to the analysis of which attention heads have strong causal influence on the LLM’s correct output. (In Mistral-7B-v0.1, each attention layer has 8 key and value activations, and 32 query activations. Therefore, heads $(\ell,h\times 4)$ to $(\ell,h\times 4+3)$ share the same key and value activation.) Patching a single head might not yield a high logit difference, since other heads in the same group (which possibly perform a similar function) could overwhelm the patched head and maintain the model’s previous “belief”. Therefore, we also run a coarser-grained experiment which simultaneously patches the attention heads sharing the same key and value activations, shown in Figure 12(b). This experiment reveals that heads belonging to the group (9, 24 - 27) also have high intervened logit difference. Combining with the observation that (9,25;26) have somewhat positive scores in the single-head patching experiments, and by examining these two head’s attention patterns (which shall be discussed in detail in the immediate next subsection), we determine that they also should be included in the circuit.
C.3.3 Attention patterns of QUERY-sensitive attention heads
@@
We can observe that head (12,9) has the “cleanest” attention pattern out of the ones identified, placing on average $93.0\pm 9.4\%$ of it attention on the correct conclusion variable alone. The more diluted attention patterns of the other heads likely contribute to their weaker intervened logit difference score shown in §4.3.1 in the main text.
@@
Extra note about head (16,0): it does not primarily act like a “mover” head, as its attention statistics suggest that it processes an almost even mixture of information from the QUERY position and the “:” position. Therefore, while we present its statistics along with the other queried-rule mover heads here since it does allocate significant attention weight on the QUERY position on average, we do not list it as such in the circuit diagram of Figure 3. Furthermore, we do not include it as part of the circuit $\mathcal{C}$ in our circuit verification experiments.
@@
Figure 15: Average attention weights of the fact processing heads computed at the “:” token position (last position before the answer), along with the standard deviations. The weights are calculated by dividing the actual attention weight placed in the Fact section by the total amount of attention placed in the relevant context (i.e. the part of the prompt excluding the 6 in-context examples).
@@
An interesting side note worth pointing out is that, (17,25) tends to only concentrate its attention in the facts section, similar to the fact-processing heads. The reason which we do not classify it as a fact-processing head and instead as a decision head is that, in addition to finding that their attention patterns tend to concentrate on the correct fact, evidence presented in §C.6 below suggest that they are not directly responsible for locating and moving the facts information to the “:” position, while the heads (16,12;14) exhibit such tendency strongly.
C.4 Sufficiency tests for circuit verification
In §4.3.2, we presented a sufficiency test of the circuit. Here, we elaborate further on the experimental procedures and finer details of the experiment.
The circuit which we perform verification on is the union of the four attention head families, $\mathcal{C}=QRLH\cup QRMH\cup FPH\cup DH$, with
[…]
  • $DH=\text{Decision Heads}=\{(19,8;9;16),(17,25)\}$ patched at the “:” position.
An exception is that the queried-rule locating head $(14,24)$ is also patched at the “:” position, as we observed that it tends to concentrate attention at the queried rule at this position: it does not locate the queried rule as consistently as it does at the QUERY position, however. We still chose to patch it at this position as we found that it tends to improve the altered logit difference, indicating that either the model relies on this head to pass certain additional information about the queried rule to the “:” position, or certain later parts of the circuit do rely on this head for queried-rule information. The exact function of this attention head remains part of our future study in the reasoning circuit of Mistral-7B. We likely need to examine this head’s role in other reasoning problems to clearly understand what its role is at different token positions, and whether there is deeper meaning behind the fact that, their apparently redundant actions at different token positions all seem to have causal influence on the model’s inference.
[…]
C.5 Queried-rule location interventions: analyzing the queried-rule locating heads
In this experiment, we only swap the location of the linear rule with the LogOp rule in the Rule section of the question, while keeping everything else the same (including all the in-context examples). As an example, we alter “Rules: A or B implies C. D implies E.” to “Rules: D implies E. A or B implies C.” while keeping everything else the same. The two prompts have the same answer.
Figure 17: Key activations patching results. In this experiment, we swap the location of the linear rule and the LogOp rule in the Rule section and keep everything else in the prompt the same; we patch the key activations of the attention heads in the Rule section only. (a) visualizes the average increase in the cross-entropy loss with respect to the true target (the true first token of the answer) for all key indices, and (b) shows the average and standard deviation of the top three key indices with the highest loss increase. Observe that these are the keys for the queried-rule locating heads (12,9), (14,24;26) and (9,25;26) identified in §4.3.1.
@@ C.5 Queried-rule location interventions: analyzing the queried-rule locating heads
The average increase in cross-entropy loss exhibit a trend which corroborate the hypothesis above, shown in Figure 17. While the average cross-entropy loss on the original samples is 0.463, patching (12,9), (14,24;26) and (9,25;26)’s keys (with corresponding key indices (12,2), (14,6) and (9,6)) in the Rule section.
Patching the other QUERY-sensitive attention heads’ keys in the Rule section, in contrast, show significantly smaller influence on the loss on average, telling us that their responsibilities are much less involved with directly finding or locating the queried rule via attention.
@@
While attention heads with key index (15,5) (i.e. heads (15, 20-23)) did not exhibit nontrivial sensitivity to QUERY-based patching (discussed in Section 4.3.1 in the main text), patching this key activation does result in a nontrivial increase in loss. Examining the attention heads belonging to this group, we find that they indeed also perform the function of locating the queried rule similar to head (12,9). We find them to be less accurate and place less attention on the exact rule being queried on average, however: this weaker “queried-rule locating ability” likely contributed to their low scores in the QUERY-based patching experiments presented in the main text.
C.6 Facts interventions: analyzing the fact-processing and decision heads
@@
C.7 Early evidence of a similar reasoning circuit in Gemma-2-9B
In this section, we present a preliminary analysis of the reasoning circuit of Gemma-2-9B in solving the same reasoning problem which Mistral-7B was examined on from before. We find that the discovered attention heads’ attention patterns inside Gemma-2-9B bear surprising resemblance to Mistral-7B’s: according to their highly specialized attention patterns, they can also be categorized into the four families of attention heads which Mistral-7B employs to solve the problem, namely the queried-rule locating heads, queried-rule mover heads, fact-processing heads, and decision heads. While it is too early to draw precise conclusions on how similar the two circuits in the two LLMs truly are, the preliminary evidence suggests that the reasoning circuit we found in this work potentially has some degree of universality.
Additionally, we caution the reader that the experimental study in this subsection is less exhaustive in nature compared to our study of Mistral-7B, due to limitations in our computation budget.
C.7.1 QUERY-based attention head activation patching
We perform activation patching of the attention head output of Gemma-2-9B, by flipping the QUERY in the prompt pairs. This is the same procedure we used to discover the attention head circuit for Mistral-7B as discussed in §4.2 and C.2. We highlight the attention heads with the strongest causal influence on the model’s (correct) inference in Figure 20.
Figure 20: QUERY-based activation patching results of Gemma-2-9B, with sub-component patching on the query and value activations. We highlight the attention heads with the highest calibrated intervened logit difference.
C.7.2 Attention patterns of QUERY-sensitive attention heads in Gemma-2-9B
Queried-rule locating heads. The queried rule locating heads inside Gemma-2-9B, namely $\{(19,11),(21,7),(22,5),(23,12)\}$, are very similar in their attention patterns to those in Mistral-7B. At the QUERY position, their attention concentrates on the conclusion token of the queried rule, and the “.” which follows. Interestingly, heads (21,7), (22,5) and (23,12) also tend to place some attention on the “implies” token of the queried rule. Another intriguing difference they exhibit is redundant behavior: these attention heads are often observed to have almost exactly the same attention pattern at the “.” and “Answer” token positions following the QUERY token. We visualize their attention statistics at the QUERY position in Figure 21.
Figure 21: Average attention weights of the queried-rule locating heads in Gemma-2-9B, along with the standard deviations. The attention pattern is obtained at the QUERY position (i.e. query activation of the attention head is from the residual stream at the QUERY token position). We record the attention weight on the queried rule.
Queried-rule mover heads. When the query activations of the queried-rule mover heads $\{(20,7),(23,6),(24,15)\}$ come from the “:” residual stream, they have fixed attention patterns which focus a large portion of their attention weights on the QUERY token and two token positions following it, namely the “.” and “Answer” token. Their attention weights are slightly more diffuse compared to their counterparts in Mistral-7B, likely due to the queried-rule locating heads performing similar functions at the “.” and “Answer” positions. Furthermore, as shown in Figure 20(c), we note that these attention heads are the only ones where patching their value activations results in a large intervened logit difference, further suggesting their role in performing a fixed “moving” action. We record their attention weights in Figure 22.
Figure 22: Average attention weights of the queried-rule mover heads in Gemma-2-9B, along with the standard deviations. The attention pattern is obtained at the “:” position, and we sum the attention weights at the QUERY position and the “.” and “Answer” token positions which immediately follow QUERY.
Fact-processing heads. The fact-processing heads $\{(24,5),(25,7),(26,0),(26,12)\}$’s attention patterns at the “:” position tend to place larger weight on the correct fact for the answer, similar to the fact-processing heads in Mistral-7B. An interesting difference does exist though: heads (24,5) and (25,7) also tend to place a nontrivial amount of weight on the QUERY and “:” token positions, indicating that these heads are relying on some form of mixture of information present at those positions for processing. While it is reasonable to hypothesize that these heads are likely relying on the queried-rule information present in the QUERY and “:” residual streams, we have not confirmed this hypothesis in our current experiments. We visualize the statistics of these heads in Figure 23.
Figure 23: Average attention weights of the fact-processing heads in Gemma-2-9B, along with the standard deviations. The attention pattern is obtained at the “:” position. We record the attention weights at the correct fact, QUERY and “:” positions, and the maximum weight on any other position.
Decision heads. The decision heads $\{(28,12),(30,9)\}$’s attention pattern are obtained at the “:” position. They bear strong resemblance to those in Mistral-7B: they place significant attention on the correct answer token (in both the rules and facts sections, same as Mistral-7B’s decision heads), and little attention weight anywhere else. This is shown in Figure 24.
Figure 24: Average attention weights of the decision in Gemma-2-9B, along with the standard deviations. The attention pattern is obtained at the “:” position. We record the attention weights at the correct answer token positions.
C.7.3 Circuit verification
We perform a sufficiency test of the attention-head circuit, following the same methodology as in the Mistral case, as discussed in §4.2 and C.4.
The circuit which we perform verification on is the union of the four attention head families, $\mathcal{C}=QRLH\cup QRMH\cup FPH\cup DH$, with
[…]
  • $DH=\text{Decision Heads}=\{(28,12),(30,9)\}$.
Remark.
The circuit verification is performed in a coarser-grained manner in this experiment, as we patch the output of the attention heads in $\mathcal{C}$ from the QUERY position to the “:” position, instead of clearly distinguishing the token positions which each head primarily focuses on.
$\mathcal{C}^{\dagger}$ $\mathcal{C}_{null}$ $\mathcal{C}$ $\mathcal{C}-QRLH$ $\mathcal{C}-QRMH$ $\mathcal{C}-FPH$ $\mathcal{C}-DH$
$\Delta^{\mathcal{C}^{\dagger}}_{orig\to alt}/\Delta_{alt}$1.0 0.94 -0.97 -0.40 0.17 -1.11
Table 2: $\Delta^{\mathcal{C}^{\dagger}}_{orig\to alt}/\Delta_{alt}$ for Gemma-2-9B, with different choices of $\mathcal{C}^{\dagger}$. $\mathcal{C}_{null}$ denotes the empty circuit, i.e. the case where no intervention is performed. We abbreviate the attention head families, for example, $DH=$ decision heads; $\mathcal{C}-DH=$ full circuit but with the decision heads removed.
We find that by patching all 13 attention heads in $\mathcal{C}$, $\Delta^{\mathcal{C}}_{orig\to alt}$ is about 94% of the “maximal” average logit difference $\Delta_{alt}$ on the altered samples. Moreover, removing any one of the four families of attention heads from $\mathcal{C}$ in the circuit interventions renders the “belief altering” effect of the intervention almost trivial.
Remark.
We find it surprising that two LLMs (Mistral-7B and Gemma-2-9B) which are trained with different procedures and data ended up relying on attention-head circuits which bear strong resemblance to each other’s. In the current literature, it is unclear how one can rigorously quantify the similarity of two nontrivial circuits inside different LLMs, however, this subsection does yield preliminary evidence that, the reasoning circuit we discover potentially has some degree of universality to it, and is likely an emergent trait of LLMs.